OpenStudy (decarr432):

Given that (-6,-5) is on the graph of f(x), find the corresponding point for the function f(3/4x)

OpenStudy (anonymous):

set \[\frac{3}{4}x=-6\]and solve

OpenStudy (decarr432):

So would it make it -24 because I have to do the opposite and multiply?

OpenStudy (triciaal):

you are on track you need x

OpenStudy (decarr432):

How could I find it? Please explain this confuses me very badly

OpenStudy (decarr432):

@triciaal

OpenStudy (anonymous):

\[\frac{3}{4}x=-6\\ x=-6\times \frac{4}{3}\] cancel l first, mult8ply last

OpenStudy (decarr432):

-8 right

OpenStudy (anonymous):

yes

OpenStudy (decarr432):

so -8 is my new x?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yes

OpenStudy (decarr432):

Thank you

OpenStudy (anonymous):

that is your new x, the y is the same

OpenStudy (anonymous):

yw

OpenStudy (decarr432):

Thank you so much help